OpenStudy (anonymous):

Factor completely: a2 – 18a + 81

OpenStudy (anonymous):

(a-9)^2

OpenStudy (anonymous):

(a-18)a+81

OpenStudy (anonymous):

=(a-9)2

OpenStudy (anonymous):

j kay your right thx

OpenStudy (gw2011):

a^2-18a+81=(a-9)(a-9)=(a-9)^2

OpenStudy (anonymous):

I had the same answer lol

OpenStudy (anonymous):

o u did sorry thanks